OpenStudy (cutestgirlaround):

13m=12–m A. 3 B. 4 C. 9 D. 12

OpenStudy (cutestgirlaround):

@Qwertty123

OpenStudy (cutestgirlaround):

I'm thinking A or B

OpenStudy (qwertty123):

Okay well... Simplify both sides of the equation. 13m=12−m

OpenStudy (qwertty123):

then Step 2: Add m to both sides. 13m+m=−m+12+m 14m=12

OpenStudy (qwertty123):

Then Divide both sides by 14. To get your answer

OpenStudy (qwertty123):

Hmm... I am not getting any of the answers from that equations that you have...

OpenStudy (cutestgirlaround):

well that would be 1.16666667 yeah I know me too

OpenStudy (qwertty123):

Nope...

OpenStudy (qwertty123):

Wait... 6/7 does that equal 1.666667

OpenStudy (cutestgirlaround):

nope it equals something like 0.857

OpenStudy (qwertty123):

oh well that is the answer I got...

OpenStudy (cutestgirlaround):

OMG I typed it in wrong its 1/3 see when I copied it from my practice test it screwed up and I for got to add the / between 1 and 3

OpenStudy (cutestgirlaround):

I'm soooooooo sorry

OpenStudy (qwertty123):

Your fine... so what is the equation?

OpenStudy (cutestgirlaround):

1/3m=12-m

OpenStudy (qwertty123):

okay then... Simplify both sides of the equation. 1 3 m=12−m

OpenStudy (qwertty123):

Add m to both sides. 1 over 3 m+m=−m+12+m 4 over 3 m=12 Multiply both sides by 3/4.

OpenStudy (qwertty123):

Then you can get your answer

OpenStudy (cutestgirlaround):

I got 9.6

OpenStudy (qwertty123):

It is 9... So C

OpenStudy (cutestgirlaround):

ok thanks again sorry for the trouble